OpenStudy (amazingmeisnotonfire):

The quadratic equation x^2+a=0 where a>0, ______ has at least one real number solution A. Always B. Sometimes C.Never

OpenStudy (amazingmeisnotonfire):

I was thinking C, is that right?

OpenStudy (phi):

if you add -a to both sides you get x^2 + a - a = 0 -a or x^2 = -a they tell you a>0, so you know -a is a negative number now take the square root x= sqr(-a) is that a real number ?

OpenStudy (amazingmeisnotonfire):

No, so is it C?

OpenStudy (phi):

yes, for a>0 you always get two imaginary answers

OpenStudy (amazingmeisnotonfire):

@phi Thank you! I've been trying to figure it out for ages